Find the inverse Laplace transform of the given function: F(S) = 3! (s – 2)
Answer #1

As we know that the Laplace transform of

L\left \{ e^{at}f(t) \right \}\Leftrightarrow F(s-a)

L\left \{ t^n\right \}\Leftrightarrow \frac{n!}{s^{n+1}}

So,

L^{-1}\left \{ F(s) \right \}=L^{-1}\left \{ \frac{3!}{(s-2)^4}\right \}

L^{-1}\left\{\frac{3!}{s^4}\right\}\Leftrightarrow \quad t^3

L^{-1}\left \{ \frac{3!}{(s-2)^4}\right \}\Leftrightarrow t^{3}e^{2t}

f(t)=e^{2t}t^3